Matthew and alex are brothers. one is 5 years older than the other. The sum of their ages are 27. what are there ages?
Answer by algebrahouse.com(1659)

x = one
x + 5 = other {one is 5 years older than the other}
x + x + 5 = 27 {sum of their ages is 27
2x + 5 = 27 {combined like terms}
2x = 22 {subtracted 5 from both sides}
x = 11 {divided both sides by 2}
x + 5 = 16 {substituted 11, in for x, into x + 5}

11 and 16 are their ages

For more help from me, visit: www.algebrahouse.com